OpenStudy (anonymous):

differentiate y= (3^x)/ (1+tan2x).

OpenStudy (anonymous):

is it \(1+\tan^2(x)\) or \(1+\tan(2x)\)?

OpenStudy (anonymous):

tan(2x) sorry I should've made that more clear haha

OpenStudy (anonymous):

ok then use \[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x)=3^x, f'(x)=\ln(3)\times 3^x, g(x)=1+\tan(2x), g'(x)=2\sec^2(2x)\] plug and chug

OpenStudy (anonymous):

thank you so much!!